OpenStudy (anonymous):

Find the value of sec theta for the angle shown. The point is (7,6)

zepdrix (zepdrix):

Hello again ms Krissie :) So we're in the first quadrant.

zepdrix (zepdrix):



zepdrix (zepdrix):

Use your Pythagorean Theorem. What is the length of the hypotenuse?

OpenStudy (anonymous):

9?

zepdrix (zepdrix):

\[\Large\rm a^2+b^2=c^2\]\[\Large\rm 7^2+6^2=c^2\]Hmm I don't think that's going to work out nicely to 9 :o

zepdrix (zepdrix):

\[\Large\rm 85=c^2\]Then square root to get your hypotenuse, yah?

zepdrix (zepdrix):



OpenStudy (anonymous):

thats what I did, i got c^2=85 and the sqrt of it is about 9

zepdrix (zepdrix):

Yah a little over 9 :) They'll probably want you to keep the answer exact though. \[\Large\rm c=\sqrt{85}\]

zepdrix (zepdrix):

So to find secant, relate it back to cosine first.\[\Large\rm \sec \theta=\frac{1}{\cos \theta}\]

zepdrix (zepdrix):

secant is the flip of cosine. Recall that:\[\Large\rm \cos \theta=\frac{adjacent}{hypotenuse}\]

zepdrix (zepdrix):

So our secant will be the flip of that,\[\Large\rm \sec \theta=\frac{hypotenuse}{adjacent}\]

OpenStudy (anonymous):

I'm so confused.. what do i plug in and where

zepdrix (zepdrix):

\[\Large\rm \sec \theta=\frac{hypotenuse}{adjacent}\]We labeled the sides. Look back at the picture :o

zepdrix (zepdrix):

\[\Large\rm \sec \theta=\frac{hypotenuse}{7}\]

zepdrix (zepdrix):

Hypotenuse was the thing we figured out using our Pythagorean Theorem, yes?

OpenStudy (anonymous):

yes, so it's 9/7 right?

OpenStudy (xapproachesinfinity):

so let's pick where @zepdrix left \(\huge {sec\theta=\frac{1}{cos\theta}} \Longrightarrow cos\theta=\frac{1}{sec\theta}=\frac{7}{\sqrt{85}}\)

OpenStudy (xapproachesinfinity):

to get sec just flip that fraction

OpenStudy (xapproachesinfinity):

@KrissieG123 the hypothenuse is not 9 is bigger than 9 just keep it \(\sqrt{85}\)

OpenStudy (anonymous):

oh okay so wait, is that the answer? or s there another step? The choices are: A. sec=(7sqrt85)/85 B. sec=6/7 C. sec=7/6 D. sec=sqrt 85/7 It would be D right?

OpenStudy (xapproachesinfinity):

yes! one you flip to get sectheta you would get the answer D

OpenStudy (anonymous):

Okay, thank you so much!!

OpenStudy (xapproachesinfinity):

np!